OpenStudy (anonymous):

If a 60-watt bulb is used to light a room, how much current is flowing through the bulb given that the voltage is 120 volts? A.180 amps B.60 amps c.2 amps d.0.5 amps

OpenStudy (anonymous):

Using P=V*I 60=120*I 60/120= .5 amps

OpenStudy (anonymous):

Answer D is the correct answer